UT 15 Sep 2021




Night Staff:

OSA: J.Williams

ISp: C. Bolyard

ISA/Observers: J.Power

Technical Wizard: J. Hill (Tucson)

Mtn Manager: P. Hartley

Sunset: 18:31

Twlight Ends: 19:49

Twilght Starts: 04:41

Sunrise: 05:59

Starting Config: LUCI-Bino (LUCI-1 authorized)

Night Log:


Proposed Plan of Action:

  • LUCI-1:
    • %Y LUCI-1 Dual BLR target - twilight until 30deg limit (Rothberg) - until 20:45 local time
      • Take Telluric first as soon as AGW-1 can wave-front sense
    • Take LUCI-1 Rotator ZPs through Astrometric Masks to confirm no changes/issues - 0.5hr
  • DONE PEPSI Training - 3 hrs
    • ready to go around 21:15 or so
    • Targets:
      • RS Oph (visible until 22:10)
      • Standard stars
    • Training Availability (priority to those using in upcoming nights)
      • Setup PEPSI and Configure as if start of night
        • Olga
        • Barry, Alex (silent mode)
  • LBC:
    • reconfigure at 00:30
    • Continue pointing model - target el > 80deg stars (run TMS in passive)
      • Skip aberration
      • First do collimation model - sort by elevation - take distribution of 10 stars covering 30 to 80ish - 1 hr
      • DONE Then do pointing model - cover sky - no collimation w/dofpia - just take short images - 2.5hr
        • 25-30 stars
        • Run Active TMS
    • If clear - standard star fields
      • DONE Prioritize Y-band scripts/fields - one field sdssY4 with active TMS
      • Can run with Active TMS (if not clear can run this anyways)
  • Other Tasks which could be done:
    • MODS:
      • Dither Test - M52 or M71 - high priority
      • Spectrophotometric stars (if clear any seeing) - all modes
      • Standard Stars (if clear any seeing)
      • Pointing Model (poor weather and poor seeing)
      • NS Longslit - low priority
    • Mixed Mode Test - if time
      • LBC-Red & MODS-1
        • Test whether LBC can send a 2nd script (Dec -90) and everything works well for both sides
    • LBC:
      • Photometric standards (any seeing - clear)
      • Spectrophotometric standards (any seeing - clear)
      • If conditions too poor for other stuff - Pointing Model in Active mode too
      • General TMS Active mode testing (any conditions)
        • run any Science-like OB in the LBC Restart OT file
        • Run for 1-1.5 hours
  • Other Poor Weather (if pointing models are exhausted)
    • LUCI-1 local mergers
      • Service_Observing/Restart_2021/LUCI/PoorWeather/COBandhead_Mergers_V17.xml - OT file
      • finder charts in sub-directory (but they are so large you can't miss 'em)

Current Status

  • LBCs
    • LBC-Blue - available
    • LBC-Red - available
  • MODS
    • MODS-1 - issues with M1RC headers & filenumbers short 1 digit
      • AGW-5 - available
    • MODS-2 - available - glycol leak - to be inspected today
      • AGW-6 - available
  • LUCI-1 - available
    • AGW-1 available
  • LUCI-2 offline
    • AGW-2 available
  • PEPSI - available
    • AGW-7/8 available


Closed Dome Activities

  • MODS-1 Full set of Cals
    • Biases
      • 1x1
        • 8k - COMPLETED
      • 1x2
      • 2x2
    • Grating
      • Dual
        • 0.8" slit, 0.3" slit, slitless pixel flats, arcs - COMPLETED
      • Red
        • 0.8" slit, slitless pixel flats, arcs - COMPLETED
      • Blue
    • Imaging
      • Dual
      • Red
      • Blue
    • Prism
      • Dual
      • Red
      • Blue
  • MODS-2 Cals
    • Bias
      • 1x2
      • 2x2
    • Prism
      • Dual
      • Red
      • Blue
  • LUCI-1 Cals
    • Imaging N375 Flats
      • Narrow, medium-band filters
      • Broad zJHKKs and HeI completed
    • N1.8 G210 CWL=2.3 K-band 0.5" cals
    • N1.8 G200 CWL=1.93 HKspec 0.5" cals
  • ARGOS Cal-Unit test for M3 alignment and pupil wobble checkout (alternate to on-sky method)
  • LUCI-1 Only Calibrations
    • Flexure data for G150
      • /home/dthomspons/Luci/FlexureCompensation
      • README - Calib15 - 2hrs
      • Ks filter, Argon lamp, optical sieve mask
      • take dark frame
      • Manual Control MCSGUI for elevations
    • Wavelength data for G150
      • G150 Ks, Argon, take image through slit, then put in grating, line should show up where the slit is located (within a few pixels of the image of the slit)
      • Check a few lines (2 or 3)

Night Overview

For the night of: Tue, 2021 Sep 14 ---> Wed, 2021 Sep 15
Local midnight = 2021 Sep 15,  7 hr UT, or JD 2459472.792
Local Mean Sidereal Time at midnight =  23 18 18.9 Sunset ( 1000 m horizon):   18 31 MST; Sunrise:   5 59 MST
Evening twilight:  19 49 MST;  LMST at evening twilight:  19 07
Morning twilight:   4 41 MST;  LMST at morning twilight:   4 00
12-degr twilight: 19 20 MST -->  5 10 MST; night center:   0 15 MST  Moonset :   0 25 MST   
Moon at civil midnight: illuminated fraction 0.653
1.4 days since first quarter, RA and dec:  18 46 58, -26 34.6  The sun is down for 11.5 hr;  8.9 hr from eve->morn 18 deg twilight.
4.3 dark hours after moonset and before beginning of twilight.

Josh has checked out the telescope. COnfigured for LUCi1 with 26 degree elevation limit. He had an SX Swing arm halting error. Mike Gardiner came on and resolved.

25 LBC biases taken.

LUCI field stop alignment verified.

1:33 Josh is opening in preparation for some LUCI sky flats.

Glenn is online in the background looking at loads on robs and looking at the rtd.

LUCI Skyflats

H2 Skyflats: luci1.20210915.0014-21

FeII Skyflats:luci1.20210915.0023-27

Ks SkyFlats: luci1.20210915.0029-38

2:04 Going for pointing and collimation near first target. Seeing is really bad. 2.1-2.8" on the guider

Dual BLR

2:12 Slewing to telluric for Dual BLR target - HIP71172
X-Offset -10.1744, Y-Offset 0.0024

2:17 Starting Telluric Science luci1.20210915.0042-43. Seeing is 1.8" on the DIMM, 1.9" on the guider.

2:21 Sending preset to Dual BLR target.

Performed mask alignment with 2.5" seeing but the seeing has just improved to 1.2-1.4". Pulling mask to turnout to verify. I'm not convinced of the alignment on cotton balls is ideal. It is difficult to discern the secondary source

Image jump during the confirmation image. This alignment is very hard to verify in variable seeing 1.4-3.5"

2:58 Science. luci1.20210915.0053-67
Seeing around 2.4" on the guider. Trace was visible in a few image subtractions, but with images of seeing <2.2 not really...

PEPSI Training

3:48 Reconfiguring for PEPSI. Seeing is too poor for astrometric mask, but we may try it later if seeing settles. Moving on to PEPSI training.

Olga is driving. First target is RS Oph.

Josh is getting copointing limits when sending PEPSI presets.

3:59 Josh is sending pointing check near our first target of the night. Struggling to find the star on the left.

2:18 Starting RS Oph. Seeing has improved. 1.4" on the DIMM. 1.3" on the guiders.

V1674 - we played with beam spliter, binnings, but we were unable to identify V1674 Her. We took a 5 min exposure and ther eare no emission lines so we are not on the right target. Back reflection? Fin

5:46 Going to Flux standard HR8634

5:50AZCAM communication lost on DX. Opened mainenance -> PFU DX ->VNC. Every looks ok. Josh restarting DX GCS. Resolved. Resending preset for DX

6:05 Flux on cross profile is just noise on DX on HR8634. Something is wrong. We will preset again. and repeat CD 1 & 6. DX collimation keep runnning away.
The cross profile again has rining and low counts. We will exposure since this is only a minute long exposure. The GCS image looks ok, although collimation could look better.

6:17 Large difference in flux from teh left or the right.Collimation looks poor on both sides. Josh is clearing the optics on both primaries
Collimation at el 68 @ 6:20UT 0.4sec WFS exp time:
Screen Shot 2021-09-14 at 11.19.22 PM.png

Collimation issues related to short exposure time on the WFS? Going to BD+28, to see if that shows similar collimation issues.

The WFS looks much, 18 sec on left and 30 on the right. Collimating well.

6:41 All done with PEPSI. Olga is setting up for PEPSI calibrations. Josh is reconfiguring for LBC's.

15/09/2021 04:18:21.3 00:00:20.000 RS Oph 5: 6278 - 7419 300 Nova 1.14 pepsir.20210915.000

15/09/2021 04:19:03.2 00:00:20.000 RS Oph 5: 6278 - 7419 300 Nova 1.11 pepsir.20210915.001

15/09/2021 04:19:44.8 00:00:40.000 RS Oph 5: 6278 - 7419 300 Nova 1.11 pepsir.20210915.002

15/09/2021 04:20:46.8 00:00:40.000 RS Oph 5: 6278 - 7419 300 Nova 1.18 pepsir.20210915.003

15/09/2021 04:21:48.8 00:00:40.000 RS Oph 5: 6278 - 7419 300 Nova 1.17 pepsir.20210915.004

15/09/2021 04:18:21.3 00:05:00.000 RS Oph 3: 4800 - 5441 300 975 Nova 1.15 pepsib.20210915.000

15/09/2021 04:22:50.3 00:00:40.000 RS Oph 5: 6278 - 7419 300 Nova 1.16 pepsir.20210915.005

15/09/2021 04:23:52.3 00:00:40.000 RS Oph 5: 6278 - 7419 300 Nova 1.10 pepsir.20210915.006

15/09/2021 04:25:05.5 00:05:00.000 RS Oph 4: 5441 - 6278 300 Nova 1.69 pepsir.20210915.007

15/09/2021 04:25:05.5 00:05:00.000 RS Oph 2: 4265 - 4800 300 144 Nova 1.69 pepsib.20210915.001

15/09/2021 04:30:55.0 00:05:00.000 RS Oph 1: 3837 - 4265 300 Nova 1.47 pepsib.20210915.002

15/09/2021 04:30:55.0 00:05:00.000 RS Oph 6: 7419 - 9067 300 Nova 1.47 pepsir.20210915.008

15/09/2021 04:39:54.6 00:02:00.000 58 Aql 5: 6278 - 7419 300 1380 Nova 1.15 pepsir.20210915.009

15/09/2021 04:39:54.6 00:02:00.000 58 Aql 3: 4800 - 5441 300 Nova 1.15 pepsib.20210915.003

15/09/2021 04:42:28.6 00:02:00.000 58 Aql 4: 5441 - 6278 300 Nova 1.01 pepsir.20210915.010

15/09/2021 04:42:28.7 00:02:00.000 58 Aql 2: 4265 - 4800 300 860 Nova 1.01 pepsib.20210915.004

15/09/2021 04:45:18.6 00:02:00.000 58 Aql 1: 3837 - 4265 300 Nova 0.96 pepsib.20210915.005

15/09/2021 04:45:18.6 00:02:00.000 58 Aql 6: 7419 - 9067 300 Nova 0.96 pepsir.20210915.011

15/09/2021 06:01:26.1 00:01:00.000 HR 8634 5: 6278 - 7419 300 fluxstd 2.15 pepsir.20210915.014

15/09/2021 06:01:26.1 00:01:00.000 HR 8634 3: 4800 - 5441 300 fluxstd 2.15 pepsib.20210915.008

15/09/2021 06:03:01.2 00:01:00.000 HR 8634 4: 5441 - 6278 300 fluxstd 1.99 pepsir.20210915.015

15/09/2021 06:03:01.2 00:01:00.000 HR 8634 2: 4265 - 4800 300 fluxstd 1.99 pepsib.20210915.009

15/09/2021 06:08:15.7 00:01:00.000 HR 8634 1: 3837 - 4265 300 fluxstd 0.94 pepsib.20210915.010

15/09/2021 06:08:15.7 00:01:00.000 HR 8634 6: 7419 - 9067 300 fluxstd 0.94 pepsir.20210915.016

15/09/2021 06:15:41.9 00:01:00.000 HR 8634 3: 4800 - 5441 200 fluxstd 1.02 pepsib.20210915.011

15/09/2021 06:15:41.9 00:01:00.000 HR 8634 5: 6278 - 7419 200 fluxstd 1.02 pepsir.20210915.017

15/09/2021 06:17:29.0 00:01:00.000 HR 8634 4: 5441 - 6278 200 fluxstd 1.00 pepsir.20210915.018

15/09/2021 06:17:29.0 00:01:00.000 HR 8634 2: 4265 - 4800 200 fluxstd 1.00 pepsib.20210915.012

15/09/2021 06:19:16.1 00:01:00.000 HR 8634 1: 3837 - 4265 200 fluxstd 0.99 pepsib.20210915.013

15/09/2021 06:19:16.1 00:01:00.000 HR 8634 6: 7419 - 9067 200 fluxstd 0.99 pepsir.20210915.019

15/09/2021 06:26:17.2 00:01:00.000 HR 8634 3: 4800 - 5441 100 fluxstd 1.16 pepsib.20210915.014

15/09/2021 06:26:17.2 00:01:00.000 HR 8634 5: 6278 - 7419 100 fluxstd 1.16 pepsir.20210915.020

15/09/2021 06:28:45.8 00:01:00.000 HR 8634 2: 4265 - 4800 100 fluxstd 1.03 pepsib.20210915.015

15/09/2021 06:28:45.8 00:01:00.000 HR 8634 4: 5441 - 6278 100 fluxstd 1.03 pepsir.20210915.021

15/09/2021 06:31:14.3 00:01:00.000 HR 8634 6: 7419 - 9067 100 fluxstd 1.08 pepsir.20210915.022

15/09/2021 06:31:14.3 00:01:00.000 HR 8634 1: 3837 - 4265 100 fluxstd 1.08 pepsib.20210915.016

7:04 Reconfiged to LBC's.

LBC Pointing Model with Active TMS

7:12 Starting dofpia on target 8, el 66 so that we can set the reference. Andrew is starting up the lasers.

Plan is to run with active tms
  • COPOINT
  • RUN 20 SEC EXP for TMS to send a few corrections
7:24 TOOK A FEW TMS DATA POINTS BEFORE SETTING THE OFFSET. SET OFFSET

SX Instr correction compensated is: 0.20055507 0.28495711 0.0 6.1134696 -4.3027081 0.0 68.51498391536092
DX Instr correction compensated is: -0.19990133 -0.25358084 0.050000001 -5.4403229 4.2886829 0.0 68.51498391536092
SX Pointing offset compensated is: -0.39384231 -0.17296827 0.0 -3.7163818 8.4620628 0.0 68.51498391536092
DX Pointing offset compensated is: -0.39383715 -0.15940553 0.0 -3.424974 8.4619522 0.0 68.51498391536092

Star az el LBCB 1 s LBCR 1s LBCB 20sec LBCR 20sec DIMM Notes
8 49 66           Initial collimation and setpoint
1 3 50     73458 73452   Forgot to turn on pointing log
                Turning on pointing log for real now
1 1 50     74857 74853   Image jump on Blue, active optics correction TMS, waiting until I see group correction from TMS before sending 20sec exposure (~15 sec). According to John this shoudl not be necessary so nevermind., Measurement taken at a bad time during copointing?
17 27 44     75436 75430    
8 36 73     80056 80054 1.1"  
24 50 34     80712 80708 0.75"  
15 54 56     81507 81502 1.12"  
25 72 31     82117 82111 0.69"  
19 87 47     82745 82739    
11 97 71     83504 83500   we have decided we have enough of the 20sec exp. Just goign straight to copointing only. taking 1 sec copointed image only after this
14 116 55 84059 85054        
16 134 33 84613 84609        
12 148 52 85305 85300     1.07"  
10 163 35 85825 85817     0.86"  
3 197 38 90422 90417       images still look great and well collimated.
5 209 77 91056 91055        
2 209 52 91646 91641        
102 237 58 92204 92200        
95 267 39 93300 93256     0.85  
100 274 55 94313 94312       offset seen in IIF but source unknown. retook image for copointing and cources look rounder.
94 296 37 95014 95009        
101 310 55 95643 95638     0.81  
8 319 72 100542 100529     1.05 Too few downloaded pixels, retaken.
93 324 36 101330 101326        
1 344 47 101859 101854     1.12  
34 368 34 102646 102642     0.88  
17 370 55 103245 103302     1.07  
15 373 76 104001 103956     1.28 Pointing quite off on this target, coma is building in blue
24               Blue images are elongated. IE and CA took a long time to send after copointing. Just hit collimation limit. Looks like I did something totally stupid. The offset I sent was huge and not reasonable. Josh is reverting the IE and CA and removing instrument offsets. Resending to 24
24             1.30 The object will visibly require a large shift for copointing. Ahhh Josh only removed instrument offsets from SX. So clearning both sides and resending. The group rotation is -3 mm. TMS is going wasky. Clearing group rotation.
24 403 65           Resent. Collimation is back in a normal range. But I still got a pretty large IE and CA for copointing... And we've hit a limit
TMS bug discovered. TMS thought the primary was ready but the primary was int a limit.
                 
We are giving up on this. There is somethign funky going on for AZ < 370

LBC Photometric Field

10:18 Slewing to sdssY4_r015d14.

10:35 new reference for tms taken after co pointing:

SX Instr correction compensated is: 0.33245304 0.47971681 0.0 10.291844 -7.1324472 0.0 57.45807754113372
DX Instr correction compensated is: -0.3340179 -0.44064206 0.025 -9.4535341 7.1660199 0.0 57.45807754113372
SX Pointing offset compensated is: -0.33765888 -0.41259217 0.0 -8.8649206 7.2549105 0.0 57.45807754113372
DX Pointing offset compensated is: -0.33764845 -0.39060846 0.0 -8.392581 7.2546864 0.0 57.45807754113372

10:37 Starting science on sdss4Y_r01514

TMS is having a regular jump in z corrsponding to filter changes.

We are noticing stray light in the upper right corner of the LBCR r images. There may be an LED on the C-ring extension? hard to tell. Mentioned it to Pat. It was not visible after we closed up.

Script failed on the last image of the script

131 2021/09/15 12:06:40.659000 CAMERA CCDCTRL safetyoff completed
130 2021/09/15 12:06:40.659000 CAMERA CCDCTRL safetyoff has been forced
129 2021/09/15 12:06:40.628000 CAMERA CCDCTRL CDS_A600_Bias_Output_Enable_by_PCI_A300 failed "FPGA not loaded" (code:48) [ccdctrl.c:2531]
128 2021/09/15 12:06:39.518000 CAMERA CCDCTRL safetyoff has been forced
127 2021/09/15 12:06:39.503000 CAMERA CCDCTRL SPC_A200_Clock_Output_Enable_by_PCI_A300 failed: "FPGA not loaded" (code:48) [ccdctrl.c:2525]

Hitting again successfully finished the last exposure. Seeing 1.18"

12:11 Script complete. Ctrl C TMS and running dopia again to get a new reference for TMS. Some coma correction in red needed.

SX Instr correction compensated is: 0.33245304 0.47971681 -0.80000001 10.291844 -7.1324472 0.0 49.20750834874714
DX Instr correction compensated is: -0.3340179 -0.44064206 -0.80000001 -9.4535341 7.1660199 0.0 49.20750834874714
SX Pointing offset compensated is: -0.35487553 -0.58970386 0.0 -12.670328 7.624826 0.0 49.20750834874714
DX Pointing offset compensated is: -0.35486558 -0.56029928 0.0 -12.038543 7.6246119 0.0 49.20750834874714

Offset taken with dofipa offset in place.

12:27 retaking offset

SX Instr correction compensated is: 0.33245304 0.47971681 0.0 10.291844 -7.1324472 0.0 47.27916851458398
DX Instr correction compensated is: -0.3340179 -0.44064206 0.025 -9.4535341 7.1660199 0.0 47.27916851458398
SX Pointing offset compensated is: -0.35487998 -0.61619776 0.0 -13.239574 7.6249213 0.0 47.27916851458398
DX Pointing offset compensated is: -0.35487029 -0.58438981 0.0 -12.55615 7.6247134 0.0 47.27916851458398

Andrew is shutting lasers off

LBC Sky Flats

halted pointing log a bit late...

12:31 Sending preset to Blank04+25 for sky flats - U & Y PA0 & PA180

Closed up

Taking 25 LBC Biases

-- BarryRothberg - 14 Sep 2021 *
Topic revision: r8 - 15 Sep 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback